My instructor made this question up. But we use the Second Edition, Intermediate Algebra testbook, by Bittinger, ellenbogen, Johnson and chapter 6 - sorry if you have answered a question like this already? I saw some similar but not with the "skill water"...
question is.
A boat travels 10 miles per hour in still water. It travels 20 miles down a river in the same time that it travels 5 miles up the same river. What is the speed of the River?
thank you.

Answer by checkley75(3666)

20/(10+X)=5/(10-X) CROSS MULTIPLY
5(10+X)=20(10-X)
50+5X=200-20X
5X+20X=200-50
25X=150
X=150/25
X=6 MPH FOR THE CURRENT
PROOF
20/16=5/4
1.25=1.25
-------------------------------------------------------------------
NEW PROBLEM: THESE ARE THE VERY EASY ONES
X*Y/(X+Y)=TOTAL TIME WHERE X & Y ARE THE TIMES IT TAKES EACH ONE INDICVIDUALLY TO DO SOMETHING OR WVWN PIPES TO FILL/EMPTY A POOL.
IN YOU PROBLEM WE HAVE A 40 MINTE JOB & AN UNKNOWN JOB TIME THAT ENDS UP BEING A 15 MINUTE JOB IF COMBINED. THUS LET X BE THE UNKNOW TIME.
40X/(40+X)=15 NOW CROSS MULTIPLY
40X=15(40+X)
40X=600+15X
40X-15X=600
25X=600 X=600/25
X=24 MINUTES FOR THE OTHER PERSON TO DO THE JOB ALONE.
PROOF
40*24/(40+24)=15
960/64=15
15=15
HAPPY MATHING